OpenStudy (anonymous):

can someone show me how to get the exact value of 3/ square root 8??

OpenStudy (jdoe0001):

no calculator around? https://desmos.com/calculator

OpenStudy (anonymous):

= 3 ⁄ √8  = 3 ⁄ √(4 • 2)  = 3 ⁄ [(√4) • (√2) ]  = 3 ⁄ (2  • √2) ... this is OK  = 1.5 ⁄ √2 ... or this  = 3  • (√2) ⁄ (2  • √2  • √2) ... multiplying by: (√2) ⁄ √2   = 3  • (√2) ⁄ 4 ... with the denominator rationalized

OpenStudy (anonymous):

8/2=4/2=2/2=1

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

Your welcome! c: